# Graph Retrieval-Augmented Generation: A Survey

### v1

@@ 6.4.2. Knowledge Enhancement
One line of methods introduces stronger models for reranking. For example, Li et al. (2023) concatenate each retrieved triplet with the question-choice pair, and adopt a pre-trained cross-encoder (Reimers and Gurevych, 2019) to re-rank the retrieved triplets. Jiang et al. (2024a) utilize the FlagEmbedding to encode the text to re-rank top-k documents returned by embedding model “bge_reranker_large”.
Another category utilizes the similarity between queries and retrieved information for ranking. For instance, Cheng et al. (2024) re-rank the candidate subgraphs based on the similarity for both relation and fine-grained concept between subgraphs and the query. Taunk et al. (2023) first cluster the 2-hop neighbors and then delete the cluster with the lowest similarity score with the input query. Yasunaga et al. (2021) prune the retrieved subgraph according to the relevance score between the question context and the KG entity nodes calculated by a pre-trained language model. Wang et al. (2024a), Jiang et al. (2023b), Gutiérrez et al. (2024) and Luo et al. (2023) adopt Personalized PageRank algorithm to rank the retrieved candidate information for further filtering. G-G-E (Gao et al., 2022) first divides the retrieved subgraph into several smaller subgraphs, then compares the similarity between each smaller subgraph and the query. Subgraphs with low similarity are removed, and the remaining smaller subgraphs are merged into a larger subgraph.
@@ 7.1.1. GNNs
Due to the powerful representational capabilities of GNNs for graph data, they are particularly effective for discriminative tasks. GNNs can directly encode graph data, capturing complex relationships and node features inherent in the graph structure. This encoding is then processed through a Multi-Layer Perceptron (MLP) to generate predictive outcomes. These approaches primarily utilize classical GNN models (e.g., GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), and Graph Transformers (Shehzad et al., 2024)), either in their original form or modified to better align with downstream tasks. For example, Sun et al. (2018) compute PageRank scores for neighboring nodes and aggregates them weighted by these scores, during message-passing. This approach enhances the central node’s ability to assimilate information from its most relevant neighboring nodes. Mavromatis and Karypis (2022) decode the query into several vectors (instructions), and enhances instruction decoding and execution for effective reasoning by emulating breadth-first search (BFS) with GNNs to improve instruction execution and using adaptive reasoning to update the instructions with KG-aware information.
@@ 7.1.3. Hybrid Models
In these methods, prompt tuning (Liu et al., 2023, 2022; Li and Liang, 2021; Lester et al., 2021) is a typical approach, where GNNs are commonly employed to encode the retrieved graph data. This encoded graph data is subsequently prepended as a prefix to the input text embeddings of an LM. The GNN is then optimized through downstream tasks to produce enhanced encodings of the graph data (He et al., 2024; Hu et al., 2024; Zhang et al., 2024b; Gao et al., 2024a).
@@
Remark. Hybrid models that harness both the representation capabilities of GNNs for graph data and LMs for text data hold promising applications. However, effectively integrating information from these two modalities remains a significant challenge.
@@ 7.2.1. Graph Languages
Remark. Good graph languages should be complete, concise, and comprehensible. Completeness entails capturing all essential information within the graph structure, ensuring no critical details are omitted. Conciseness refers to the necessity of keeping textual descriptions brief to avoid the “lost in the middle” phenomenon (Liu et al., 2024a) or exceeding the length limitations of LMs. Lengthy inputs can hinder LMs’ processing capabilities, potentially causing loss of context or truncated data interpretation. Comprehensibility ensures that the language used is easily understood by LLMs, facilitating accurate representation of the graph’s structure. Due to the characteristics of different graph languages, their choice can significantly impact the performance of downstream tasks (Fatemi et al., 2023).
@@ 7.2.2. Graph Embeddings
The above graph language methods transform graph data into text sequences, which may result in overly lengthy contexts, incurring high computational costs and potentially exceeding the processing limits of LLMs. Additionally, LLMs currently struggle to fully comprehend graph structures even with graph languages (Guo et al., 2023). Thus, using GNNs to represent graphs as embeddings presents a promising alternative. The core challenge lies in integrating graph embeddings with textual representations into a unified semantic space. Current research focuses on utilizing prompt tuning methodologies, as discussed earlier. Notably, feeding graph representations into LMs is feasible primarily with open-source LMs, not closed-source models like GPT-4 (OpenAI, 2024). While graph embedding methods avoid handling long text inputs, they face other challenges, such as difficulty in preserving precise information like specific entity names and poor generalization.
@@ 7.3.3. Post-Generation Enhancement
Post-generation enhancement occurs after the initial response is generated. Post-generation enhancement methods primarily involve integrating multiple generated responses to obtain the final response. Some methods focus on integrating outputs from the same generator under different conditions or inputs. For example, Edge et al. (2024) generate a summary for each graph community, followed by generating responses to queries based on the summary, and then scoring these responses using an LLM. Ultimately, the responses are sorted in descending order according to their scores and sequentially incorporated into the prompt until the token limit is reached. Subsequently, the LLM generates the final response. Wang et al. (2023c) and Kim et al. (2023a) first decompose the query into several sub-questions, then generate answers for each sub-question, and finally merge the answers of all sub-questions to obtain the final answer.
Alternatively, other methods combine or select responses generated by different models. Lin et al. (2019) and Jiang et al. (2022) combine the outputs generated by both GNNs and LLMs to reach a synergistic effect. UniOQA (Li et al., 2024a) explores two methods for generating answers: one involves generating queries in Cypher Query Language (CQL) to execute and obtain results, while the other method directly generates answers based on retrieved triplets. The final answer is determined through a dynamic selection mechanism. In EmbedKGQA (Saxena et al., 2020), besides the learned scoring function, researchers additionally design a rule-based score based on the graph structures. These two scores are combined to find the answer entity. Li et al. (2024f) combine answers based on retrieved graph data with responses generated according to the LLM’s own knowledge.
@@ 8.1.2. Training-Based
Training retrievers often adopt an autoregressive approach, where the previous relationship path is concatenated to the end of the query. The model then predicts the next relation based on this concatenated input (Wu et al., 2023b; Guo et al., 2024).
However, the lack of ground truth for retrieval content in the majority of datasets poses a significant challenge. To address this issue, many methods attempt to construct reasoning paths based on distant supervision to guide retriever training. For example, Zhang et al. (2022b), Feng et al. (2023) and Luo et al. (2024b) extract all paths (or shortest paths) between entities in the queries and entities in the answers, using them as training data for the retriever. In addition, Zhang et al. (2022b) also employ a relationship extraction dataset for distant supervision in unsupervised settings. There is another category of methods that utilize implicit intermediate supervision signals to train Retrievers. For instance, KnowGPT (Zhang et al., 2024a) starts searching for the optimal path from the head entity, using the discovery of the tail entity as a reward, and is trained using Policy Gradient. NSM (He et al., 2021) employs a bidirectional search strategy, where two retrievers start searching from the head entity and tail entity, respectively. The supervised objective is to ensure that the paths searched by the two retrievers converge as closely as possible.
Some methods argue that distant supervision signals or implicit intermediate supervision signals may contain considerable noise, making it challenging to train effective retrievers. Therefore, they consider employing self-supervised methods for pre-training retrievers. SKP (Dong et al., 2023) pretrains the DPR (Dense Passage Retrieval) model (Karpukhin et al., 2020). Initially, it conducts random sampling on subgraphs and transforms the sampled subgraphs into passages. Subsequently, it randomly masks passages, trains the model using a Masked Language Model (MLM), and employs contrastive learning by treating the masked passages and original passages as positive pairs for comparison.
@@ 9. Applications and Evaluation
QA KBQA WebQSP (Yih et al., 2016) (Luo et al., 2024b), (Sun et al., 2024b), (Zhang et al., 2022b), (Wu et al., 2023b), (Guo et al., 2024), (Wang et al., 2023a), (Jiang et al., 2023a), (Jiang et al., 2024b), (Luo et al., 2024a), (Wang et al., 2023c), (Baek et al., 2023b), (Shu et al., 2022), (Liu et al., 2024d), (Mavromatis and Karypis, 2024), (Sun et al., 2019), (Yu et al., 2023), (Dong et al., 2023), (Gao et al., 2022), (Luo et al., 2023), (Baek et al., 2023a), (Saxena et al., 2020), (He et al., 2021), (Jiang et al., 2023b), (Choi et al., 2023), (Sun et al., 2018), (Jiang et al., 2023c) Accuracy, Hits@1, EM, Recall, F1, BERTScore, GPT-4 Average Ranking
WebQ (Berant et al., 2013) (Wang et al., 2024a)(Sun et al., 2024b)(Hu et al., 2022)(Wu et al., 2023b)(Mavromatis and Karypis, 2022)(Jiang et al., 2023c)
CWQ (Talmor and Berant, 2018) (Luo et al., 2024b), (Sun et al., 2024b), (Hu et al., 2022), (Zhang et al., 2022b), (Wang et al., 2023a), (Jiang et al., 2024b), (Luo et al., 2024a), (Mavromatis and Karypis, 2022), (Lan and Jiang, 2020), (Liu et al., 2024d), (Mavromatis and Karypis, 2024), (Sun et al., 2019), (Yu et al., 2023), (Gao et al., 2022), (Luo et al., 2023), (Jiang et al., 2023b), (Li et al., 2024f), (He et al., 2021), (Choi et al., 2023)
GrailQA (Gu et al., 2021) (Sun et al., 2024b)(Jiang et al., 2024b)(Shu et al., 2022)
QALD10-en (Perevalov et al., 2022) (Sun et al., 2024b)(Li et al., 2024f)(Sun et al., 2024a)
SimpleQuestions (Bordes et al., 2015) (Sun et al., 2024b)(Baek et al., 2023b)
@@
MetaQA (Zhang et al., 2018) (Luo et al., 2024b)(Wu et al., 2023b)(Guo et al., 2024)(Kim et al., 2023a)(Wang et al., 2023c)(Mavromatis and Karypis, 2022)(Sun et al., 2019)(Saxena et al., 2020)(Jiang et al., 2023a)(Liu et al., 2024d)(He et al., 2021)(Jiang et al., 2023b)(Choi et al., 2023)
Natural Question (Kwiatkowski et al., 2019) (Hu et al., 2022)
@@
HotpotQA (Yang et al., 2018) (Hu et al., 2022)(Gutiérrez et al., 2024)
FACTKG (Kim et al., 2023b) (Kim et al., 2023a)
Mintaka (Sen et al., 2022) (Baek et al., 2023b)(Li et al., 2024f)(Baek et al., 2023a)
@@
CSQA CSQA (Talmor et al., 2019) (Taunk et al., 2023)(Yasunaga et al., 2021)(Huang et al., 2023a)(Li et al., 2023)(Lin et al., 2019)(Feng et al., 2023)
OBQA (Mihaylov et al., 2018) (Taunk et al., 2023)(Yasunaga et al., 2021)(Huang et al., 2023a)(Li et al., 2023)(Feng et al., 2023)(Han et al., 2023)
@@
ZsRE (Petroni et al., 2021) (Li et al., 2024f)(Sun et al., 2024a)(Sun et al., 2024b)
Others Fact Verification Creak (Onoe et al., 2021) (Li et al., 2024f)(Sun et al., 2024a)(Sun et al., 2024b) Hits@1
@@
Recommender Systems Yelp [^9] https://www.yelp.com/dataset/ (Wang et al., 2022) NDCG@$K$, Recall@$K$
@@ 9.1.3. Others
In addition to the aforementioned downstream tasks, GraphRAG can be applied to various other tasks in the realm of natural language processing such as fact verification, link prediction, dialogue system, and recommender systems.
@@
(4) Recommender Systems
In the context of e-commerce platforms, the purchase relationships between users and products naturally form a network graph. The primary objective of recommender systems within these platforms is to predict the future purchasing intentions of users, effectively forecasting the potential connections within this graph (Wang et al., 2022).

### v2

@@ 6.4.2. Knowledge Enhancement
One line of methods introduces stronger models for reranking. For example,  Li et al. (2023) concatenate each retrieved triplet with the question-choice pair, and adopt a pre-trained cross-encoder (Reimers and Gurevych, 2019) to re-rank the retrieved triplets.  Jiang et al. (2024a) utilize the FlagEmbedding to encode the text to re-rank top-k documents returned by embedding model “bge_reranker_large”.  Liu et al. (2024b) train a PLM to
Another category utilizes the similarity between queries and retrieved information for ranking. For instance,  Cheng et al. (2024) re-rank the candidate subgraphs based on the similarity for both relation and fine-grained concept between subgraphs and the query.  Taunk et al. (2023) first cluster the 2-hop neighbors and then delete the cluster with the lowest similarity score with the input query.  Yasunaga et al. (2021) prune the retrieved subgraph according to the relevance score between the question context and the KG entity nodes calculated by a pre-trained language model.  Wang et al. (2024a) , Jiang et al. (2023b) , Gutiérrez et al. (2024) and Luo et al. (2023) adopt Personalized PageRank algorithm to rank the retrieved candidate information for further filtering.  Liu et al. (2024b) trains a PLM to score the similarity between the retrieved information and the query, and rerank the retrieved paths based on the similarity score. G-G-E (Gao et al., 2022) first divides the retrieved subgraph into several smaller subgraphs, then compares the similarity between each smaller subgraph and the query. Subgraphs with low similarity are removed, and the remaining smaller subgraphs are merged into a larger subgraph.
@@ 7.1.1. GNNs
Due to the powerful representational capabilities of GNNs for graph data, they are particularly effective for discriminative tasks. GNNs can directly encode graph data, capturing complex relationships and node features inherent in the graph structure. This encoding is then processed through a Multi-Layer Perceptron (MLP) to generate predictive outcomes. These approaches primarily utilize classical GNN models (e.g., GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), and Graph Transformers (Shehzad et al., 2024)), either in their original form or modified to better align with downstream tasks. For example, HamQA (Dong et al., 2023b) designs a hyperbolic GNN to learn the representations of retrieved graph data, which learns from the mutual hierarchical information between query and graphs.  Sun et al. (2018) compute PageRank scores for neighboring nodes and aggregates them weighted by these scores, during message-passing. This approach enhances the central node’s ability to assimilate information from its most relevant neighboring nodes.  Mavromatis and Karypis (2022) decode the query into several vectors (instructions), and enhances instruction decoding and execution for effective reasoning by emulating breadth-first search (BFS) with GNNs to improve instruction execution and using adaptive reasoning to update the instructions with KG-aware information.
@@ 7.1.3. Hybrid Models
In these methods, prompt tuning (Liu et al., 2023, 2022; Li and Liang, 2021; Lester et al., 2021) is a typical approach, where GNNs are commonly employed to encode the retrieved graph data. The encoded graph data is subsequently pre-pended as a prefix to the input text embeddings of an LM. The GNN is then optimized through downstream tasks to produce enhanced encodings of the graph data (He et al., 2024; Hu et al., 2024; Zhang et al., 2024b; Gao et al., 2024a).
@@
Discussion
Hybrid models that harness both the representation capabilities of GNNs for graph data and LMs for text data hold promising applications. However, effectively integrating information from these two modalities remains a significant challenge.
@@ 7.2.1. Graph Languages
Discussion
Good graph languages should be complete, concise, and comprehensible. Completeness entails capturing all essential information within the graph structure, ensuring no critical details are omitted. Conciseness refers to the necessity of keeping textual descriptions brief to avoid the “lost in the middle” phenomenon (Liu et al., 2024a) or exceeding the length limitations of LMs. Lengthy inputs can hinder LMs’ processing capabilities, potentially causing loss of context or truncated data interpretation. Comprehensibility ensures that the language used is easily understood by LLMs, facilitating accurate representation of the graph’s structure. Due to the characteristics of different graph languages, their choice can significantly impact the performance of downstream tasks (Fatemi et al., 2023).
@@ 7.2.2. Graph Embeddings
The above graph language methods transform graph data into text sequences, which may result in overly lengthy contexts, incurring high computational costs and potentially exceeding the processing limits of LLMs. Additionally, LLMs currently struggle to fully comprehend graph structures even with graph languages (Guo et al., 2023). Thus, using GNNs to represent graphs as embeddings presents a promising alternative. The core challenge lies in integrating graph embeddings with textual representations into a unified semantic space. Current research focuses on utilizing prompt tuning methodologies, as discussed earlier. There are also some methods that adopt FiD (Fusion-in-Decoder) (Izacard and Grave, 2021; Yu et al., 2022), which first convert the graph data into text, then encode it using an LM-based encoder and input it into the decoders (Fang et al., 2024a; Yu et al., 2023; Dong et al., 2023a). Notably, feeding graph representations into LMs is feasible primarily with open-source LMs, not closed-source models like GPT-4 (OpenAI, 2024). While graph embedding methods avoid handling long text inputs, they face other challenges, such as difficulty in preserving precise information like specific entity names and poor generalization.
@@ 7.3.3. Post-Generation Enhancement
Post-generation enhancement occurs after the initial response is generated. Post-generation enhancement methods primarily involve integrating multiple generated responses to obtain the final response. Some methods focus on integrating outputs from the same generator under different conditions or inputs. For example,  Edge et al. (2024) generate a summary for each graph community, followed by generating responses to queries based on the summary, and then scoring these responses using an LLM. Ultimately, the responses are sorted in descending order according to their scores and sequentially incorporated into the prompt until the token limit is reached. Subsequently, the LLM generates the final response.  Wang et al. (2023c) and Kim et al. (2023a) first decompose the query into several sub-questions, then generate answers for each sub-question, and finally merge the answers of all sub-questions to obtain the final answer. Alternatively, other methods combine or select responses generated by different models.  Lin et al. (2019) and Jiang et al. (2022) combine the outputs generated by both GNNs and LLMs to reach a synergistic effect. UniOQA (Li et al., 2024a) explores two methods for generating answers: one involves generating queries in Cypher Query Language (CQL) to execute and obtain results, while the other method directly generates answers based on retrieved triplets. The final answer is determined through a dynamic selection mechanism. In EmbedKGQA (Saxena et al., 2020), besides the learned scoring function, researchers additionally design a rule-based score based on the graph structures. These two scores are combined to find the answer entity.  Li et al. (2024f) combine answers based on retrieved graph data with responses generated according to the LLM’s own knowledge. In addition to integrating multiple responses, KALMV (Baek et al., 2023c) trains a verifier to judge whether the generated answer is correct, and if it is not, to further determine whether the error is due to generation or retrieval.
@@ 8.1.2. Training-Based
When the retrieval granularity is nodes or triplets, many methods train retrievers to maximize the similarity between the retrieval ground truth and the query. For instance, MemNNs (Bordes et al., 2015a) leverages metric learning to closely align the ground truth with the query in semantic space while differentiating unrelated facts from the query. On the contrary, when the retrieval granularity is paths, training retrievers often adopts an autoregressive approach, where the previous relationship path is concatenated to the end of the query. The model then predicts the next relation based on the concatenated input (Wu et al., 2023b; Guo et al., 2024).
However, the lack of ground truth for retrieval content in the majority of datasets poses a significant challenge. To address this issue, many methods attempt to construct reasoning paths based on distant supervision to guide retriever training. For example,  Zhang et al. (2022b) , Feng et al. (2023) and Luo et al. (2024b) extract all paths (or shortest paths) between entities in the queries and entities in the answers, using them as training data for the retriever. In addition,  Zhang et al. (2022b) also employ a relationship extraction dataset for distant supervision in unsupervised settings. There is another category of methods that utilize implicit intermediate supervision signals to train Retrievers. For instance, NSM (He et al., 2021) employs a bidirectional search strategy, where two retrievers start searching from the head entity and tail entity, respectively. The supervised objective is to ensure that the paths searched by the two retrievers converge as closely as possible. KnowGPT (Zhang et al., 2024a) and MINERVA (Das et al., 2018) treat the selection of adjacent nodes to build paths or subgraphs as a Markov process. They design the reward function around the inclusion of the answer in the retrieved information and adopt reinforcement learning methods e.g. policy gradient to optimize the retriever.
Some methods argue that distant supervision signals or implicit intermediate supervision signals may contain considerable noise, making it challenging to train effective retrievers. Therefore, they consider employing self-supervised methods for pre-training retrievers. SKP (Dong et al., 2023a) pre-trains the DPR (Dense Passage Retrieval) model (Karpukhin et al., 2020). Initially, it conducts random sampling on subgraphs and transforms the sampled subgraphs into passages. Subsequently, it randomly masks passages, trains the model using a Masked Language Model (MLM), and employs contrastive learning by treating the masked passages and original passages as positive pairs for comparison.
@@ 9. Applications and Evaluation
QA KBQA WebQSP (Yih et al., 2016) (Luo et al., 2024b), (Sun et al., 2024b), (Ma et al., 2024), (Zhang et al., 2022b), (Wu et al., 2023b), (Guo et al., 2024), (Wang et al., 2023a), (Jiang et al., 2023a), (Jiang et al., 2024b), (Luo et al., 2024a), (Wang et al., 2023c), (Baek et al., 2023b), (Shu et al., 2022), (Liu et al., 2024e), (Mavromatis and Karypis, 2024), (Sun et al., 2019), (Yu et al., 2023), (Dong et al., 2023a), (Gao et al., 2022), (Luo et al., 2023), (Baek et al., 2023a), (Saxena et al., 2020), (He et al., 2021), (Jiang et al., 2023b), (Choi et al., 2023), (Sun et al., 2018), (Jiang et al., 2023c), (Baek et al., 2023c), (Dehghan et al., 2024) Accuracy, EM, Recall, F1, BERTScore, GPT-4 Average Ranking
WebQ (Berant et al., 2013) (Wang et al., 2024a)(Sun et al., 2024b)(Hu et al., 2022)(Wu et al., 2023b)(Mavromatis and Karypis, 2022)(Jiang et al., 2023c)(Bordes et al., 2015a)(Dehghan et al., 2024)
CWQ (Talmor and Berant, 2018) (Luo et al., 2024b), (Sun et al., 2024b), (Hu et al., 2022), (Zhang et al., 2022b), (Wang et al., 2023a), (Jiang et al., 2024b), (Luo et al., 2024a), (Mavromatis and Karypis, 2022), (Lan and Jiang, 2020), (Liu et al., 2024e), (Mavromatis and Karypis, 2024), (Sun et al., 2019), (Yu et al., 2023), (Gao et al., 2022), (Luo et al., 2023), (Jiang et al., 2023b), (Li et al., 2024f), (He et al., 2021), (Choi et al., 2023), (Dehghan et al., 2024)
GrailQA (Gu et al., 2021) (Sun et al., 2024b)(Jiang et al., 2024b)(Shu et al., 2022)(Dehghan et al., 2024)
QALD10-en (Perevalov et al., 2022) (Sun et al., 2024b)(Ma et al., 2024)(Li et al., 2024f)(Sun et al., 2024a)
SimpleQuestions (Bordes et al., 2015b) (Sun et al., 2024b)(Baek et al., 2023b)(Bordes et al., 2015a)(Dehghan et al., 2024)
@@
MetaQA (Zhang et al., 2018) (Luo et al., 2024b)(Wu et al., 2023b)(Guo et al., 2024)(Kim et al., 2023a)(Wang et al., 2023c)(Mavromatis and Karypis, 2022)(Sun et al., 2019)(Saxena et al., 2020)(Jiang et al., 2023a)(Liu et al., 2024e)(He et al., 2021)(Jiang et al., 2023b)(Choi et al., 2023)(Liu et al., 2024b)
Natural Question (Kwiatkowski et al., 2019) (Hu et al., 2022)(Baek et al., 2023c)(Dehghan et al., 2024)
@@
HotpotQA (Yang et al., 2018) (Hu et al., 2022)(Gutiérrez et al., 2024)(Ma et al., 2024)(Baek et al., 2023c)(Dehghan et al., 2024)
Mintaka (Sen et al., 2022) (Baek et al., 2023b)(Li et al., 2024f)(Baek et al., 2023a)(Baek et al., 2023c)
@@
CSQA CSQA (Talmor et al., 2019) (Taunk et al., 2023)(Yasunaga et al., 2021)(Huang et al., 2023a)(Li et al., 2023)(Lin et al., 2019)(Feng et al., 2023)(Dong et al., 2023b)
OBQA (Mihaylov et al., 2018) (Taunk et al., 2023)(Yasunaga et al., 2021)(Huang et al., 2023a)(Li et al., 2023)(Feng et al., 2023)(Han et al., 2023)(Dong et al., 2023b)
@@
ZsRE (Petroni et al., 2021) (Li et al., 2024f)(Sun et al., 2024a)(Sun et al., 2024b)(Ma et al., 2024)
Others Fact Verification Creak (Onoe et al., 2021) (Li et al., 2024f)(Sun et al., 2024a)(Sun et al., 2024b)(Ma et al., 2024) Accuracy, F1
FACTKG (Kim et al., 2023b) (Kim et al., 2023a)(Lan and Jiang, 2020)(Liu et al., 2024b)
@@
Recommendation Yelp [^9] https://www.yelp.com/dataset/ (Wang et al., 2022) NDCG@$K$, Recall@$K$
@@ 9.1.3. Others
In addition to the aforementioned downstream tasks, GraphRAG can be applied to various other tasks in the realm of natural language processing such as fact verification, link prediction, dialogue systems, and recommendation.
@@
(4) Recommendation
In the context of E-commerce platforms, the purchase relationships between users and products naturally form a network graph. The primary objective of recommendation within these platforms is to predict the future purchasing intentions of users, effectively forecasting the potential connections within this graph (Wang et al., 2022).